30 Apr (Thursday)

  • Operations dominated by MC jobs and user analysis
  • T0:
    • VOMS server retirement on Tuesday, all ok
  • T1

27 Apr (Monday)

  • Operations dominated by MC jobs and user analysis
  • T0:
    • Problem with SRM for pit file export (GGUS:113316)
    • Discussing old VOMS servers retirements, previously planned for Wed, might happen on Tue now
  • T1

23 Apr (Thursday)

  • Operations dominated by MC jobs and user analysis
  • VOMS-Admin ticket GGUS:112282 currently on hold
  • few issues with T2 sites, NIPNE FTS transfer failures being looked into (GGUS:112044), pilots to DESY-ZN new CEs at first not being served, fixed now (GGUS:112700)

20 Apr (Monday)

  • Operations dominated by MC jobs and user analysis
  • VOMS-Admin ticket GGUS:112282 currently on hold

16 April (Thursday)

  • Mostly user and MC jobs in the system.
  • FTS transfer problems to Nipne continuing (GGUS:112044)
  • VOMS-Admin issues waiting for solution from VOMS admins : GGUS:112282 and GGUS:112279

13 April (Monday)

  • User and MC jobs are running in the system. The failure rate of the jobs is very low.
  • CERN & Tier-1s running fine
    • Problem of failing LHCb jobs solved promptly by RAL admins last Friday evening

09 April (Thursday)

  • User and MC jobs are running in the system. The failure rate of the jobs is very low.

  • No issues to report

-- JoelClosier - 2015-06-01

Edit | Attach | Watch | Print version | History: r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r1 - 2015-06-01 - JoelClosier
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LHCb All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback